OpenStudy (anonymous):

What is the d value of the following arithmetic sequence? -16, -9, -2, 5, 12, 19….

OpenStudy (anonymous):

7

OpenStudy (anonymous):

is -9 -(-16)=-9+16=7

mathslover (mathslover):

d = a_2 - a_1 d= a_3 -a_2

OpenStudy (anonymous):

Is it really that difficult? kids these days (sigh..)

mathslover (mathslover):

-9 - (-16) = -9 + 16 = 7 -2 -(-9) = -2+9 = 7

mathslover (mathslover):

hence d = 7 got it @coolaidd ?

mathslover (mathslover):

"no medals yet" quite ridicolous : D

mathslover (mathslover):

*ridiculous

OpenStudy (anonymous):

thank you!

mathslover (mathslover):

np